OpenStudy (anonymous):

Hi guys, I am trying to find the derivative using l'Hospital's Rule. The problem is: lim x-> pi/2 ==> (1-sin)/(1+cos(2x)) The answer is: 1/4 My trig skills aren't very good. Can someone help me answer this?

OpenStudy (anonymous):

Basically with L'Hopital's rule you use the derivative of the function to find the limit (that's not a very good defn actually but it is essentially what you are doing) For this question you want\[\lim_{x \rightarrow \frac{ \pi }{ 2 }} \frac{ 1-\sin(x) }{ 1+\cos(2x) }\] if you take the derivative, you get: \[\lim_{x \rightarrow \frac{ \pi }{ 2 }}\frac{ 1 }{ 4\sin(x) }\] If you factor out the constants, as x approaches pi/2, sin(x) tends to 1 So the whole limit will tend to 1/4 Does that make sense?

OpenStudy (anonymous):

can u tell me how you got 4sin(x)

OpenStudy (anonymous):

because i understand that cos(2x) = -2*sin(2x) right?

OpenStudy (psymon):

Well, just take the derivaitve of it. It's a chain rule that would giveyou -2sin(2x)

OpenStudy (anonymous):

omg

OpenStudy (psymon):